Pablo wants to build a pen for the pig he is raising in his agriculture class. Pablo only has 36 feet of fencing. What dimensions can Pablo use to give the pen the maximum area? Organize your information into a 4-corner model.
then it asks for
Scatterplot/Graph:
Dimensions for maximum are and verbal description:
Domain and range:

Answer by richwmiller(17219)

A square would give the maximum area.
4s=p
9*4=36
Each side is 9 feet
Area=9^2=81